OpenStudy (anonymous):

-4x+4y=-7 5x-6y=7 I cant eliminate here,right? So what do i do?

OpenStudy (anonymous):

What do you mean?

OpenStudy (slaaibak):

First multiply each of the equations by a different factor to let either the x coefficients or y coefficients be the same. Them just add the two so the one variable would be cancelled out.

OpenStudy (turingtest):

You can eliminate them as slaaibak said. think lowest common multiple between the coefficients of one of the variables.

OpenStudy (anonymous):

Oh i see, so i could do this? -4x+4y=-7-------------12x+12y=-21 5x-6y=7--------------10x-12y=14 To cancel out the 12s?

OpenStudy (turingtest):

bingo

OpenStudy (anonymous):

So i would have -12x+12y=-21 + 10x-12y=14 = -2x=7?

OpenStudy (turingtest):

yep

OpenStudy (anonymous):

-2x=-7

OpenStudy (turingtest):

sorry, -7 careful oh you caught it

OpenStudy (anonymous):

-2x=-7 =-7/2?

OpenStudy (turingtest):

yep yep! and now for y...

OpenStudy (turingtest):

no, sorry positive again

OpenStudy (turingtest):

negative over negative...

OpenStudy (anonymous):

Substitute 5x-6y=7 getting 5*7/2-6y=7

OpenStudy (anonymous):

Oh right

OpenStudy (turingtest):

you caught it anyway so far so good, you're doing better than me this morning lol

OpenStudy (anonymous):

35/10-6y=7?

OpenStudy (anonymous):

Lol xd

OpenStudy (turingtest):

where did the ten come from?

OpenStudy (anonymous):

Oh, so i dont multiply 5 by both parts of the factor? 35/2-6y=7?

OpenStudy (turingtest):

much better :)

OpenStudy (turingtest):

in general we have this situation:\[c*\frac{a}{b}=\frac{c}{1}*\frac{a}{b}=\frac{ca}{b}\]so when a number is alone we always treat it as over one...

OpenStudy (anonymous):

Ohhhh, now i get it!

OpenStudy (turingtest):

try to relate it to those rules for algebra so what is y?

OpenStudy (anonymous):

35/2-6y=7 Im not really sure what to do here, maybe make the two sides factors match. 35/2-6y=14/2?

OpenStudy (turingtest):

that's a good start now get y on one side and the constants on another

OpenStudy (anonymous):

35/2-6y=14/2 -6y=-21/2

OpenStudy (anonymous):

Good so far?

OpenStudy (turingtest):

yes :)

OpenStudy (anonymous):

-6y=-21/2 -y=-(21/2)/6

OpenStudy (anonymous):

No that cant be it

OpenStudy (anonymous):

I need to make the factor divisible by six

OpenStudy (turingtest):

first off take the -1 with the 6, we don't want y to be negative.. second it is right, but you need to simplify it remember that dividing by a number is the same as multiplying by it's inverse

OpenStudy (anonymous):

That would mean that i have to divide both sides by -6, right?

OpenStudy (turingtest):

right, and what is (-21/2)/(-6) ??? do you know the trick fro this?

OpenStudy (anonymous):

y=-42/4 / -6? y=7/4

OpenStudy (turingtest):

yes!!! dang you learned a lot Inopeki, nice :D many people here hate systems, it'll be nice to have you help them :P

OpenStudy (anonymous):

:D Yes!!

OpenStudy (turingtest):

that's very impressive doing systems like that, I can honestly say I'm impressed. Again, awesome :D

OpenStudy (anonymous):

Thanks :D

OpenStudy (anonymous):

Give me another one :D

OpenStudy (turingtest):

well, systems can be tricky to make up off the top of your head. If you're not careful they can get messy...

OpenStudy (turingtest):

ok here's one... x-2y=4 3x+6y=3 I think this is a bit easier than the last...

OpenStudy (anonymous):

im sorry, OS crashed

OpenStudy (turingtest):

I'm in no rush, don't worry

OpenStudy (anonymous):

x-2y=4-----------3x-6y=12 3x+6y=3----------3x+6y=3

OpenStudy (anonymous):

pellet

OpenStudy (anonymous):

They will cancel out, i cant do that

OpenStudy (anonymous):

no wait

OpenStudy (anonymous):

3x-6y=12 + 3x+6y=3 = 6x=15 15/6=x?

OpenStudy (turingtest):

yep, just simplify the fraction

OpenStudy (anonymous):

5/2?

OpenStudy (turingtest):

sweet :) and for y...?

OpenStudy (anonymous):

5/2-2y=8/2 -2y=13/2?

OpenStudy (turingtest):

watch the signs... made a little mistake

OpenStudy (anonymous):

Ah, -2y=3/2?

OpenStudy (turingtest):

there ya go !

OpenStudy (anonymous):

-2y=3/2 y=6/4 / -2 y=-3/4?

OpenStudy (turingtest):

very nice! do you know how to check your answers?

OpenStudy (anonymous):

Yeah, put the final answers in to one of the equations to see if it is true. 5/2-2*3/4=4

OpenStudy (anonymous):

Now that just doesnt make sense with fractions...

OpenStudy (turingtest):

you should really put it in both equations, but yes you got the idea

OpenStudy (turingtest):

it is correct: 5/2-2(-3/4)=5/2+3/2=8/2=4 again, you dropped a negative for the y-value. That happens to me a lot in systems too, just gotta be careful

OpenStudy (turingtest):

try to verify that the other equation is true

OpenStudy (anonymous):

The first one in this question?

OpenStudy (turingtest):

the other equation in the system: 3x+6y=3

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

3*5/2+6*-3/4=3 15/2- -18/4=3 15/2- -9/2=6/2?

OpenStudy (anonymous):

=3

OpenStudy (turingtest):

so it must be right :) nicely done

OpenStudy (anonymous):

Thanks :D Now what?

OpenStudy (turingtest):

you should always check both equations, I'm not sure who told you that checking just one is okay, it's really not. Watch: x-2y=4 3x+6y=3 no what if I got the answer x=6, y=1 the first equation would give 6-2(1)=4 True but the second... 3(6)+6(1)=3 False so you must check both each equation by itself has infinite solutions, you need to use the other equation to just get one.

OpenStudy (turingtest):

now what if*

OpenStudy (anonymous):

Oh, now i see why.

OpenStudy (turingtest):

because if we solve say the first one for x... x=2y+4 we have a function again: y=1 x=2(1)+4=6 y=2 x=2(2)+4=8 etc... next? I guess systems with 2 unknowns when you're comfortable with this stuff. What else are you doing? quadratics?

OpenStudy (anonymous):

After this, its an introduction to ratios

OpenStudy (anonymous):

Oh you mean what i HAVE been doing?

OpenStudy (turingtest):

I meant the next step would be with 3 unknowns, you already are doing 2... but yeah, let's practice things you have just learned recently

OpenStudy (anonymous):

I leanred averages,taking procent, slope, y-intercept,graphing lines, graphing points, integer sums.

OpenStudy (anonymous):

equation of a line, equation systems

OpenStudy (turingtest):

lol that's a lot... what do you mean by integer sums?

OpenStudy (anonymous):

Like if the sum of 5 odd consecutive integers is 217, what is the largest odd integer?

OpenStudy (anonymous):

Actually thats 7 odd consecutive integers.

OpenStudy (turingtest):

nice those are fun... well for a lot of us at least. what do you feel unsure about? let's work on that but I need to get breakfast...

OpenStudy (anonymous):

Yeah, i need to get something to eat too

OpenStudy (turingtest):

ok we'll regroup in 10-15

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

Why does it say that im level 23?

OpenStudy (turingtest):

ok I'm back now what to work on... let's see how you are with distribution real quick: simplify 2x(3+4x-2/x)

OpenStudy (anonymous):

2x(3+4x-2/x) 2(3+4x-2) 2(1+4x)

OpenStudy (turingtest):

careful, that 3 is lonely...

OpenStudy (anonymous):

What?

OpenStudy (turingtest):

and x(4x) is not 4x

OpenStudy (anonymous):

wait!

OpenStudy (turingtest):

middle step ...

OpenStudy (anonymous):

2x^2(3-2)?

OpenStudy (turingtest):

not quite. Middle step: 2x(3+4x-2/x)=2x(3)+2x(4x)-2x(2/x)=??? ^^^^^very important step for you, you'll see why when you do factoring

OpenStudy (anonymous):

Oh right!